Write all the factors of the following numbers: 
21 ?????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

21 = 1 × 21 21 = 3 × 7 21 = 7 × 3
Stop here since 3 and 7 have occurred earlier
Hence, the factors of 21 are 1, 3, 7 and 21

Thanks